OpenStudy (hellokitty17):

What are the factors of x2 - 49?

OpenStudy (anonymous):

Do you Have any answer choices ?

OpenStudy (anonymous):

Remember earlier how I said \[x^2-49 = (x-7)(x+7)\] take each of those results let them = 0 and solve for x, woo!

OpenStudy (hellokitty17):

What are the factors of x2 - 49? (x - 1)(x + 49) (x - 7)(x - 7) (x - 7)(x + 7) Prime

oregonduck (oregonduck):

x^2 - 49 is known as a "difference of two perfect squares". (You'll see a lot of those in math classes.) Memorize the equation: a^2 - b^2 = (a + b)(a - b) x^2 - 49 = (x + 7)(x - 7)

OpenStudy (anonymous):

The factors are already given!